# Data-adaptive structural change-point detection via isolation

### v1

@@ 5 Simulations
Method $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 1 99 0 0 0 0.262 0.292 0.016
ID_th 0 0 0 69 26 5 0 0.269 0.317 0.013
ID_ic 0 0 0 97 3 0 0 0.201 0.245 0.396
CPOP 0 0 0 97 3 0 0 0.388 0.235 0.445
NOT 100 0 0 0 0 0 0 4.730 99.000 0.075
MARS 100 0 0 0 0 0 0 4.700 98.500 0.004
TF 0 0 0 0 0 0 100 0.205 0.387 0.617
TS 0 36 32 32 0 0 0 0.858 0.579 1.380
Table 12: Distribution of $\hat{N}-N$ of 100 simulated data sequences of (S10). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 0 100 0 0 0 0.038 0.203 0.016
ID_th 0 0 0 86 14 0 0 0.042 0.270 0.011
ID_ic 1 4 17 78 0 0 0 0.106 1.720 0.353
CPOP 0 0 0 98 2 0 0 0.252 0.294 0.072
NOT 100 0 0 0 0 0 0 1.060 119.000 0.048
MARS 100 0 0 0 0 0 0 1.060 118.000 0.003
TF 0 0 0 0 0 3 97 0.190 0.324 0.299
TS 100 0 0 0 0 0 0 1.060 119.000 0.478
@@ 6.1 Crime data
In this section, DAIS is applied to real data and the underlying signal is assumed to be piecewise-constant. The chosen dataset includes daily crime reports and can be found in https://catalog.data.gov/dataset/crime. We use daily observations starting from 22/01/2020, up to 31/07/2023. Since the data involve counts of crimes, they are positive integer-valued and thus a transformation is required to bring them closer to Gaussian data with constant variance. This is done using the Anscombe transform, [3], $\alpha:\mathbb{N}\rightarrow\mathbb{R}$, with $\alpha(x)=2\sqrt{x+3/8}$.
The top plot of Figure 3 is a plot of the true data, before the transformation, along with the estimated underlying signal, $\hat{f}_{t}$, according to the change-points detected by DAIS, plotted in red. The fits obtained by three other algorithms, ID, MOSUM and NOT, are also shown in the bottom plot of the figure, plotted with blue, yellow and green lines, respectively. All methods obtain different signals on the same data.
[…]
[scale=1]Tikz_Plots/crime_tikz_plot
Figure 3: The observed time series data of daily crime reports in Montgomery County and fitted piecewise-constant mean signal obtained by applying DAIS (red), ID (blue), MOSUM (yellow) and NOT (green) to the data.
@@
We now attempt to provide a possible explanation about the change-points estimated by DAIS that express the most important movements in the data. The first change-point occurs on the $17^{\textrm{th}}$ of March 2020 and corresponds to the first days of positive cases of COVID-19 when the first rules were imposed to limit the spread of the virus among the community in the state of Maryland, where Montgomery County is located. As expected, since people were forced to stay at home, the number of crimes reported dropped significantly. The second change-point is on the $28^{\textrm{th}}$ of April 2020, when the official authorities started lifting the restrictions and the number of crimes increased. Some small decrease around the $11^{\textrm{th}}$ of December 2020 can be explained by people’s behavioral change around the holidays. The number of crimes returns close to the previous levels, on the $13^{\textrm{th}}$ of May 2021 and a further increase is observed at the end of the summer, $19^{\textrm{th}}$ of August. This could possibly be due to people going on holidays. As has also been observed for 2020, there is again a reduced number of reported crimes around the Christmas holidays of 2021, starting from the $23^{\textrm{th}}$ of December 2021 until the $17^{\textrm{th}}$ of January 2022. It is notable that the next change-point detected is almost a year later, as there is a sudden decrease on the $23^{\textrm{rd}}$ of December 2022, with an increase on the $26^{\textrm{th}}$.
It is important to note that there is a clear trend as change-points were detected by DAIS on 11/12/2020, 23/12/2021 and 23/12/2022. All three are near the Christmas holidays and a decrease in the daily number of crimes reported is noticed, followed by the trend increasing on the 13/05/2021, 17/01/2022 and 26/12/2023. ID detects the same movements in the data, detecting all 6 change-points mentioned above, while NOT only misses the change on 13/05/2021. MOSUM detects a change-point close to Christmas of 2020 only. Interestingly, when the same data were used up to 15/06/2023, ID and NOT were only able to detect the change-points around Christmas 2020 and 2021. Using less data DAIS was able to detect the same change-points that two well-established methods managed to identify only after the addition of more information. It seems that the performance of DAIS is at least as good as that of some of the best-performing methods available in the literature.
@@ 6.2 Euro to British pound exchange rate
The DAIS algorithm for changes in the slope can be applied to the Euro to British pound exchange rate. The data are weekly Close prices adjusted for splits for the period between 07/04/2014 and 11/08/2023. The results can be seen in Figure 4. The first plot is the observed signal, while the other plots are the estimated signals from DAIS, ID and CPOP. All four plots have the same range of values on the y-axis so that they are comparable.
All three fits are very similar, with ID detecting the least number of change-points (25) and thus having a smoother signal, while DAIS and CPOP detect 41 and 57, respectively. From the plots it can be seen that around the start of 2016, there is an increase in the value of the exchange rate. This could be explained by the fact that 2016 was the year of the referendum where Brexit was voted by the citizens. It is worth noting that the drop in March of 2020, that is only detected by DAIS, coincides with the day that the Prime Minister Boris Johnson announced the Government would be implementing measures intended to halt the spread of the COVID-19 virus. The big drop around June of 2022 falls around the time when fuel prices soared, according to https://www.theguardian.com/theguardian/2022/jun/09.
[…]
[scale=0.8]Tikz_Plots/Real_data_slope
Figure 4: The observed time series data of Euro to British pound sterling exchange rate and fitted piecewise-linear signal obtained by applying DAIS (red), ID (blue) and CPOP (green).
@@ 7 Conclusions
In this project, DAIS is introduced, which is a new data-adaptive method for detecting structural changes in a given data sequence. The first step of the method is to identify the location of the largest difference in the sequence, as defined in (2) for the cases of changes in the mean and changes in the slope, in an effort to start searching from an area around the most apparent change-point in the case that there is one. Expanding intervals around the location of the largest difference are used to achieve isolation of the change-points, which enhances the detection power. The theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations are given in Section 3. The simulation results presented in Section 5, indicate that DAIS is at least as accurate as the state-of-the-art competitors, with its data-adaptive nature being advantageous regarding the detection of the true change-points in difficult structures, for example when the change-points are really close to each other. DAIS uses intervals around the location of the largest difference and so the change-points are, with high probability, near the midpoint of these intervals, which results in the power of the contrast function being maximized. As explained in Sections 2.3 and 3, for the proof of the theoretical properties of DAIS we require that the location of the largest difference is at a location where detection, while the change-point is isolated in an interval, is guaranteed. However, for the correct choice of the parameters, the points that are problematic are just the midpoints between two consecutive change-points. Finally, in practice, the algorithm performs at least as well as other widely-used algorithms in the literature.
Appendix A Simulation Study
@@
Appendix B Proof of Theorem 1

### v2

@@ 6.2 Heavy-tailed noise
$d$ Signal $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (ms)
(S1) 0 1 0 87 6 6 0 0.025 0.038 1.0
5 (S7) 0 0 0 70 14 12 4 0.016 0.081 1.9
(S8) 0 0 0 37 22 21 20 2 21.1 10
(S1) 0 0 0 92 4 3 1 0.022 0.023 1.0
7 (S7) 0 0 0 83 6 8 3 0.015 0.060 0.8
(S8) 0 0 0 82 8 10 0 2.01 21 10
6.3 Multivariate models
The DAIS algorithm can be extended to multivariate or high dimensional models. The model considered is the following:
$\boldsymbol{X_{t}}=\boldsymbol{f_{t}}+\boldsymbol{\epsilon_{t}},t=1,...,T,$ (20)
where $\boldsymbol{X_{t}}\in\mathbb{R}^{d\times 1}$ are the observed data and $\boldsymbol{f_{t}}\in\mathbb{R}^{d\times 1}$ is the d-dimensional deterministic signal with structural changes at certain points. In the case of piecewise-constant signals, the structure of $\boldsymbol{f_{t}}$ is given by $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j}}$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}}}\neq\boldsymbol{f_{r_{j}+1}}$ where $\boldsymbol{\mu_{j}}\in\mathbb{R}^{d\times 1}$ for $j=1,2,\ldots,N+1$. For continuous piecewise-linear signals, $\boldsymbol{f_{t}}=\boldsymbol{\mu_{j,1}}+\boldsymbol{\mu_{j,2}}t$ for $t\in\{r_{j-1}+1,\ldots,r_{j}\}$ and $\boldsymbol{f_{r_{j}-1}}+\boldsymbol{f_{r_{j}+1}}\neq 2\boldsymbol{f_{r_{j}}}$ where $\boldsymbol{\mu_{j,1}},\boldsymbol{\mu_{j,2}}\in\mathbb{R}^{d\times 1}$. To ensure continuity, we need the additional constraint of $\boldsymbol{\mu_{k,1}}+\boldsymbol{\mu_{k,2}}r_{k}=\boldsymbol{\mu_{k+1,1}}+\boldsymbol{\mu_{k+1,2}}r_{k}$ for $k\in\{1,2,\ldots,N\}$. As with the univariate case, the algorithm can also be applied to more general signal structures.
The calculation of the largest difference, that was used as a starting point for the univariate signal, needs to be adapted for the multivariate case. This can be done using mean-dominant norms $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, whose definition can be found in [9]. Some examples are
$L_{2}(\boldsymbol{X_{t}}):=\sqrt{\frac{1}{d}\sum_{i=1}^{d}X_{t,i}^{2}},\quad L_{\infty}(\boldsymbol{X_{t}}):=\sup_{i=1,\ldots,d}\lvert X_{t,i}\rvert.$ (21)
The location of the largest difference of the interval $[s,e]$, $1\leq s<e\leq T$ can be calculated as
$d^{\text{multi}}_{s,e}=\left\{\begin{array}[]{ll}\textrm{argmax}_{t\in\{s,s+1,\dots,e-1\}}\left\{L(\lvert\boldsymbol{X_{t+1}}-\boldsymbol{X_{t}}\rvert)\right\},&\boldsymbol{f_{t}}\textrm{ piecewise-constant},\\ \textrm{argmax}_{t\in\{s,s+1,\dots,e-2\}}\left\{L(\lvert\boldsymbol{X_{t+2}}-2\boldsymbol{X_{t+1}}+\boldsymbol{X_{t}}\rvert)\right\},&\boldsymbol{f_{t}}\textrm{ piecewise-linear.}\end{array}\right.$ (22)
We define $\boldsymbol{y_{t}}$ to be the $\mathbb{R}^{d\times 1}$ vector with entries the value of the chosen contrast function, this being (10) for piecewise-constant signals and (12) for piecewise-linear ones. At each step of the algorithm, the goal is to decide whether the point that maximizes $L(\boldsymbol{y_{t}})$ is a change-point. The aggregation scheme for the contrast function, as just described, was also proposed by [3]. Algorithm 2 provides a pseudocode for DAIS in the case of multivariate signals, which we call MDAIS. For a discussion on the choice of the mean-dominant norm, we refer the reader to [1, 3, 12, 15].
Theoretical results for this variant of the algorithm about consistency for the number and locations of the estimated change-points can be proven. In the rest of this section, we provide the consistency theorem that concerns the case of a piecewise-constant $\boldsymbol{f_{t}}$ and the outline of its proof can be found in Appendix E. However, since the focus of this paper is proposing the novel data-adaptive methodology, we do not delve further in this. Following a similar methodology, theoretical results can also be proved for $\boldsymbol{f_{t}}$ being continuous piecewise-linear. For the statement of the theorem, we require the following assumption, which is equivalent to assumption (A2) for univariate $f_{t}$:
  1. (A4)
    The quantities $\delta_{T}$ and $\underline{f}_{T}$ are connected by $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}_{M}\sqrt{\log\left(Td^{1/4}\right)}$, for a large enough constant $\underline{C}_{M}$.
Similarly to the univariate case, the number of change-points $N$ is allowed to grow with $T$ and $d$. Note that the threshold constant, $\zeta_{T,d}$ depends on both the length of the data sequence and its dimension.
Theorem 3.
Let $\{\boldsymbol{X_{t}}\}_{t=1,\ldots,T}$ follow model (20) with $\boldsymbol{f_{t}}$ piecewise-constant and let $\boldsymbol{\epsilon_{t}}\sim\mathcal{N}_{d}(\boldsymbol{0},\Sigma)$, where $\Sigma\in\mathbb{R}^{d\times d}$ is positive definite and also that assumptions (A1) and (A4) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ their estimates, sorted in increasing order. In addition, $\boldsymbol{\Delta_{j}}:=\lvert\boldsymbol{f_{r_{j}+1}}-\boldsymbol{f_{r_{j}}}\rvert,j=1,2,\ldots N$ is the magnitude of each discontinuity in $\boldsymbol{f_{t}}$, $\underline{f}_{T}:=\inf_{j=1,\ldots,N}\{L\left(\boldsymbol{\Delta_{j}}\right)\}$ and $\delta_{T}=\min_{j=1,2,\ldots,N+1}\lvert r_{j}-r_{j-1}\rvert$, where $r_{0}=0,r_{N+1}=T$. Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ which do not depend on $T$ such that for $C_{1}\sqrt{\log\left(Td^{1/4}\right)}\leq\zeta_{T,d}<C_{2}\sqrt{\delta_{T}}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
$\mathbb{P}\Biggl{(}\hat{N}=N,\max_{j=1,2,\ldots,N}\biggl{\{}\left|\hat{r}_{j}-r_{j}\right|\alpha^{2}\biggl{\}}\leq C_{3}\log\left(Td^{1/4}\right)\Biggl{)}\geq 1-\frac{C_{4}}{T}.$ (23)
where $\alpha=\Delta^{q_{j}}_{j}=\lvert f_{r_{j+1},q_{j}}-f_{r_{j},q_{j}}\rvert$ when $L(\cdot)=L_{\infty}(\cdot)$, and $\alpha=L_{2}\left(\boldsymbol{\Delta_{j}}\right)$ when $L(\cdot)=L_{2}(\cdot)$, $q_{j}:=\textrm{argmax}_{k=1,\ldots,d}\tilde{X}^{\hat{r}_{j},k}_{s_{j},e_{j}}$ for $[s_{j},e_{j}]$ being the interval where $\hat{r}_{j}$ is obtained and $\tilde{X}_{s,e}^{b,j}=\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum_{t=s}^{b}X_{t,j}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum_{t=b+1}^{e}X_{t,j}$, $\ell=e-s+1$.
The outline of the proof of Theorem 3 is given in Appendix E. Note that this extension of our algorithm is robust to spatial dependence. Theoretical results related to this statement, as well as simulation results can be found in [3], where the extension of ID ([2]), namely the MID algorithm, is introduced.
Algorithm 2 MDAIS
function MDAIS$\left(s,e,\lambda_{T},\zeta_{T},L(\cdot)\right)$
if $e-s<1$ then
     STOP
else
     Set $d^{\text{multi}}_{s,e}$ as in (22)
     For $j\in\{1,2,\dots,K\}$ let $I_{j}=[s_{j},e_{j}]$ as in (4)
     $i=1$
     (Main part)
     $b_{i}={\rm{argmax}}_{t\in[s_{i},e_{i})}L\left(C_{s_{i},e_{i}}^{t}\left(\boldsymbol{X}\right)\right)$
     if $L\left(C_{s_{i},e_{i}}^{b_{i}}\left(\boldsymbol{X}\right)\right)>\zeta_{T}$ then
         add $b_{i}$ to the list of estimated change-points
         MDAIS($s,s_{i},\lambda_{T},\zeta_{T}$)
         MDAIS($e_{i},e,\lambda_{T},\zeta_{T}$)
     else
         $i=i+1$
         if $i\leq K$ then
              Go back to (Main part) and repeat
         else
              STOP
         end if
     end if
end if
end function
@@ 7.1 Crime data
In this section, DAIS is applied to real data and the underlying signal is assumed to be piecewise-constant. The chosen dataset includes daily crime reports in Montgomery County, Maryland and can be found in https://catalog.data.gov/dataset/crime. We use daily observations starting from 22/01/2020, up to 31/08/2024. Since the data involve counts of crimes, they are positive integer-valued and thus a transformation is required to bring them closer to Gaussian data with constant variance. This is done using the Anscombe transform, [4], $\alpha:\mathbb{N}\rightarrow\mathbb{R}$, with $\alpha(x)=2\sqrt{x+3/8}$.
@@
The top plot of Figure 3 is a plot of the real data (in black), along with the estimated underlying signal, $\hat{f}_{t}$, according to the change-points detected by DAIS (plotted in red). We now attempt to provide a possible explanation about the change-points estimated by DAIS that express the most important movements in the data. The first change-point occurs on the $17^{\textrm{th}}$ of March 2020 and corresponds to the first days of positive cases of COVID-19 when the first rules were imposed to limit the spread of the virus among the community in the state of Maryland. As expected, since people were forced to stay at home, the number of crimes reported dropped significantly. The second change-point is on the $28^{\textrm{th}}$ of April 2020, when the official authorities started lifting the restrictions and the number of crimes increased. Some small decrease around the $11^{\textrm{th}}$ of December 2020 can be explained by people’s behavioural change around the holidays. The number of crimes returns close to the previous levels, on the $28^{\textrm{th}}$ of May 2021 and a further increase is observed at the end of the summer, $27^{\textrm{th}}$ of August 2021. This could possibly be due to people going on holidays out of town, leaving their homes more vulnerable to break-ins. As has also been observed for 2020, there is again a reduced number of reported crimes around the Christmas holidays of 2021, starting from the $23^{\textrm{th}}$ of December 2021 until the $17^{\textrm{th}}$ of January 2022. It is notable that the next change-point detected is almost a year later, as there is a sudden decrease on the $23^{\textrm{rd}}$ of December 2022, with an increase on the $26^{\textrm{th}}$ of the same month. An increase in the daily number of crimes reported is also detected on the $28^{\text{th}}$ of September 2023. As in the previous years, there is a drop on the $23^{\textrm{rd}}$ of December 2023 and an increase to a slightly higher average than before on the $23^{\textrm{rd}}$ of January 2024. Further increase to the number of daily crimes reported is observed on the $26^{\textrm{th}}$ of February and the $10^{\textrm{th}}$ of April 2024, with a decrease on the $4^{\textrm{th}}$ of June 2024. It is clear that the average number of crimes has increased in the last years. It is important to note that there is a clear trend as change-points were detected by DAIS on 11/12/2020, 23/12/2021, 23/12/2022 and 23/12/2023. All four are near the Christmas holidays and a decrease in the daily number of crimes reported is noticed, followed by the trend increasing on the 17/01/2022, 26/12/2022 and 22/01/2024.
The fits obtained by three competitors, ID, MOSUM, and NOT, are shown in the bottom plot of Figure 3, plotted with blue, yellow and green lines, respectively. All methods obtain different signals on the same data, with ID and DAIS providing similar fits and NOT having minor differences with them. ID and NOT detect 11 and 10 out of the 15 change-points that DAIS detects, respectively. They both don’t detect the change-points in May 2021 and the last 2 change-points detected by DAIS, in April and June of 2024. ID also does not detect the change-point just before Christmas 2020, while NOT does not detect the 2 change-points around Christmas 2022. MOSUM detects only 5 change-points, with the only change-point around the Christmas period being the one in 2020, which ID missed. By detecting change-points around the same period every year, we could argue that DAIS is detecting the seasonality around the Christmas holidays, which none of the other competitors managed to fully capture.
Figure 3: The observed time series data of daily crime reports in Montgomery County and fitted piecewise-constant mean signal obtained by applying DAIS (red), ID (blue), MOSUM (yellow) and NOT (green) to the data. The x-axis labels correspond to the month and year.
@@ 7.2 Euro to British pound exchange rate
The DAIS algorithm for changes in the slope is applied to the Euro to British pound exchange rate. The data are the weekly close prices adjusted for splits for the period between 07/04/2014 and 12/09/2024. The results can be seen in Figure 4. The first plot is the observed signal, while the other plots are the estimated piecewise-linear signals from DAIS, ID and CPOP. The change-points which are discussed in the next paragraph are indicated by dashed lines in the plot of the observed signal.
All three fits are very similar, with ID detecting the least number of change-points (27) and thus having a smoother signal, while DAIS and CPOP detect 42 and 60, respectively. The first important change-point, detected only by DAIS, on the week of $5^{\textrm{th}}$ of January of 2015 is the same week as the Charlie Hebdo shooting, the terrorist attack that occurred at the offices of the satirical weekly newspaper in Paris. After this, the exchange rate decreased, indicating that the attack had negative effects to the strength of the Euro. The next important change-point is an increase in the value of the exchange rate starting on the week of $9^{\textrm{th}}$ of November 2015. This change-point is detected by both DAIS and CPOP, while ID detects one just one week later. The change might have been caused by Prime Minister David Cameron’s speech on the $10^{\textrm{th}}$ of November, during which he repeated his commitment to holding a referendum for Brexit before the end of 2017. A few days later, on the $13^{\textrm{th}}$, the terrorist attacks in Paris took place. Both events had an impact on the exchange rate. The next change-point, from which a new increase in the observed value begins, is detected in the week of $13^{\textrm{th}}$ of June 2016, which is the week just before the referendum was held, on the $23^{\textrm{rd}}$ of June. It is worth noting that the drop in the week of the $16^{\text{th}}$ of March of 2020, that is detected by DAIS and CPOP, coincides with the day that the Prime Minister Boris Johnson announced the government would be implementing measures intended to halt the spread of the COVID-19 virus. The big drop around June of 2022, with change-points detected by all methods in two consecutive weeks, on the $6^{\text{th}}$ and $13^{\text{th}}$, falls around the time when fuel prices soared, according to https://www.theguardian.com/theguardian/2022/jun/09.
Figure 4: The observed time series data of Euro to British pound sterling exchange rate and fitted piecewise-linear signal obtained by applying DAIS (red), ID (green) and CPOP (blue).
@@ 8 Conclusions
In this project, DAIS is introduced, which is a new data-adaptive method for detecting structural changes in a given data sequence. The first step of the method is to identify the location of the largest difference in the sequence, as defined in (2) for the cases of changes in the mean and changes in the slope, in an effort to start searching from an area around the most apparent change-point in the case that there is one. This step is what makes our method data-adaptive and differentiates it from the competitors. Expanding intervals around the location of the largest difference, in absolute value, are used to achieve isolation of the change-points, which enhances the detection power. Theoretical results regarding the consistency of the number of change-points detected and the accuracy of their estimated locations are provided. The simulation results presented indicate that DAIS is at least as accurate as the state-of-the-art competitors, with its data-adaptive nature being advantageous regarding the detection of the true change-points in difficult structures, for example when the change-points are really close to each other. DAIS uses intervals around the location of the largest difference and so the change-points are, with high probability, near the midpoint of these intervals, which results in the power of the contrast function being maximized. For the proof of the theoretical properties of DAIS we require that the location of the largest difference is at a location where detection, while the change-point is isolated in an interval, is guaranteed. However, for the correct choice of the parameters, the points that are problematic are just the midpoints between two consecutive change-points. In practice, the algorithm performs at least as well as other widely-used algorithms in the literature, exhibiting uniformly good behaviour across different signal structures. Finally, some extensions to more difficult signal structures are discussed, which include heavy tailed noise, temporal dependence and the extension of the method to multivariate signals.
Appendix A Signals used in the simulation study
@@
    $many\_cpts\_mix$: piecewise-constant signal of length 75 with 11 change-points at 5, 12, 17, 25, 31, 38, 44, 50, 56, 61, 67 with values between the change-points 0, 5, 0, 6, 0, 4, 0, 5, 0, 6, 0, 4. The standard deviation is $\sigma=1$.
  • (S5)
@@
  • (S6)
    $many\_cpts\_long$: piecewise constant signal of length 600 with 119 change-points at 5, 10, 15, $\ldots$, 595 with values between change-points 0, 5, 0, $\ldots$, 0, 5. The standard deviation is $\sigma=1$.
  • (S7)
    $simple\_signal$: sequence of length 1100 with 1 change-point at 550 with values between change-points 0, 2. The standard deviation is $\sigma=1$.
@@
Appendix B Further simulations
Some further simulation results are presented in Tables 8 and 9 for piecewise-constant signals and Tables 10 and 11 for continuous piecewise-linear signals. The signals used are:
  • (S11)
  • (S12)
  • (S13)
  • (S14)
  • (S15)
  • (S16)
  • (S17)
    $wave5$: continuous piecewise-linear signal of length 350 with 50 change-points at 7, 14, $\ldots$, 343 with changes in slope -2.5, 2.5, $\ldots$, 2.5, -2.5. The starting intercept is $f_{1}=-0$ and slope $f_{2}-f_{1}=1$. The standard deviation is $\sigma=1$.
The general conclusions form the tables are that DAIS performs at least as well as the competitors, often having an advantage in computational time and accuracy compared to the best performing competitors. Signal (S11), in Table 8 has no change-points. Most algorithms have very good results with negligible differences in the accuracy in the number of change-points detected, WBS_th with the threshold constant $C=\sqrt{2}$ seems to overestimate. Similar conclusions hold for signal (S12), in which the big length of the signal forces some algorithms, especially MSCP, to be extremely slow. Moreover, comparing DAIS with ID_th which use the same value for the expansion parameter $\lambda_{T}$, it is evident that DAIS requires a smaller number of expansions and thus is faster. Signal (S13) in Table 9 indicates that in this difficult structure, where pairs of change-points are close and move in opposite directions such that they ‘cancel’ each-other out, DAIS and MOSUM have the best performance, in terms of accuracy in the number of change-points detected and MSE. For the signal (S14) all algorithms perform similarly, with DAIS being one of the fastest methods. The low computational complexity of DAIS can also be seen in Table 11.
Table 8: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S11) and (S12). The average MSE and computational times are also given.
$\hat{N}-N$
Method Signal -1 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 0 1 0 $1.87\times 10^{-4}$ 0.237
ID_th 0 84 8 7 1 $5.26\times 10^{-4}$ 0.450
ID_ic 0 99 1 0 0 $1.68\times 10^{-4}$ 0.080
WBS_th 0 52 26 20 2 $7.00\times 10^{-4}$ 0.067
WBS_ic 0 100 0 0 0 $1.39\times 10^{-4}$ 0.066
WBS2 (S11) 0 91 5 3 1 $2.75\times 10^{-4}$ 3.51
PELT 0 100 0 0 0 $1.39\times 10^{-4}$ 0.003
NOT 0 100 0 0 0 $1.39\times 10^{-4}$ 0.066
MOSUM 0 94 2 4 0 $2.96\times 10^{-4}$ 0.028
MSCP 0 100 0 0 0 $1.39\times 10^{-4}$ 83.25
SeedBS_th 0 12 23 25 40 $12.9\times 10^{-4}$ 0.027
SeedBS_ic 0 99 1 0 0 $1.68\times 10^{-4}$ 0.027
DP_univar 0 0 0 0 100 $372\times 10^{-4}$ 30.70
DAIS 0 99 0 1 0 $4.33\times 10^{-4}$ $43.10\times 10^{-4}$ 0.322
ID_th 0 88 1 10 1 $7.77\times 10^{-4}$ $35.80\times 10^{-4}$ 0.416
ID_ic 0 100 0 0 0 $4.25\times 10^{-4}$ $1.09\times 10^{-4}$ 0.110
WBS_th 0 75 19 6 0 $5.15\times 10^{-4}$ $626.00\times 10^{-4}$ 0.107
WBS_ic 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.107
WBS2 (S12) 0 92 3 4 1 $5.46\times 10^{-4}$ $17.60\times 10^{-4}$ 5.950
PELT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.004
NOT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.103
MOSUM 0 92 2 6 0 $5.49\times 10^{-4}$ $203.60\times 10^{-4}$ 0.049
MSCP 6 93 1 0 0 $341.35\times 10^{-4}$ $334.74\times 10^{-4}$ 276.294
SeedBS_th 0 10 23 35 42 $10.2\times 10^{-4}$ $3140\times 10^{-4}$ 0.0418
SeedBS_ic 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.0418
DP_univar 0 0 0 0 100 $368\times 10^{-4}$ $4840\times 10^{-4}$ 190.80
Table 9: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S13) and (S14). The average MSE, $d_{H}$ and computational times are also given.
$\hat{N}-N$
Method Signal $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 15 3 80 0 2 0 0.0321 0.067 0.004
ID_th 0 2 2 75 8 7 6 0.0323 0.037 0.003
ID_ic 2 34 1 62 1 0 0 0.0350 0.143 0.002
WBS_th 0 0 1 31 28 26 14 0.0336 0.088 0.025
WBS_ic 2 33 1 64 0 0 0 0.0337 0.140 0.025
WBS2 (S13) 0 6 5 70 9 6 4 0.0306 0.132 0.479
PELT 26 65 0 9 0 0 0 0.0596 0.437 0.001
NOT 2 30 1 67 0 0 0 0.0333 0.128 0.024
MOSUM 0 13 2 79 3 3 0 0.0287 0.063 0.015
MSCP 95 5 0 0 0 0 0 0.1390 0.593 2.610
SeedBS_th 0 0 1 18 27 14 40 0.0381 0.097 0.013
SeedBS_ic 1 29 2 63 3 1 1 0.0331 0.122 0.013
DP_univar 0 0 0 2 6 4 88 0.0602 0.126 0.139
DAIS 0 0 1 94 4 1 0 2.72 ×$\times 10^{-2}$ 0.010 0.002
ID_th 0 0 1 84 12 2 1 2.72 ×$\times 10^{-2}$ 0.011 0.001
ID_ic 0 0 0 87 12 1 0 2.46 ×$\times 10^{-2}$ 0.009 0.002
WBS_th 0 0 0 56 21 19 4 2.58 ×$\times 10^{-2}$ 0.016 0.008
WBS_ic 0 0 0 93 7 0 0 2.35 ×$\times 10^{-2}$ 0.008 0.008
WBS2 (S14) 0 0 1 96 3 0 0 2.31 ×$\times 10^{-2}$ 0.008 0.088
PELT 1 6 3 90 0 0 0 2.63 ×$\times 10^{-2}$ 0.016 0.001
NOT 0 0 0 95 5 0 0 2.47 ×$\times 10^{-2}$ 0.009 0.021
MOSUM 0 0 0 95 3 2 0 2.26 ×$\times 10^{-2}$ 0.008 0.008
MSCP 100 0 0 0 0 0 0 21.10 ×$\times 10^{-2}$ 0.261 0.242
SeedBS_th 0 0 1 49 25 15 10 2.68 ×$\times 10^{-2}$ 0.017 0.011
SeedBS_ic 0 0 0 93 5 2 0 2.40 ×$\times 10^{-2}$ 0.090 0.012
DP_univar 49 18 2 31 0 0 0 8.89 ×$\times 10^{-2}$ 0.210 0.003
Table 10: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signal (S15). The average MSE and computational times are also given.
$\hat{N}-N$
Method Signal 0 1 2 $\geq 3$ MSE Time (s)
DAIS 100 0 0 0 $2.12\times 10^{-3}$ 0.032
ID_th 83 10 4 3 $2.45\times 10^{-3}$ 0.066
ID_ic 100 0 0 0 $2.12\times 10^{-3}$ 0.007
CPOP 100 0 0 0 $2.45\times 10^{-3}$ 3.220
NOT (S15) 100 0 0 0 $2.31\times 10^{-3}$ 0.062
MARS 0 100 0 0 $5.25\times 10^{-3}$ 0.006
TF 7 44 37 12 $5.29\times 10^{-3}$ 0.429
TS 100 0 0 0 $2.12\times 10^{-3}$ 0.618
local_poly 10 5 17 68 $6.47\times 10^{-3}$ 33.000
Table 11: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S16) and (S17). The average MSE, $d_{H}$ and computational times are also given.
$\hat{N}-N$
Method Signal $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 0 96 4 0 0 0.022 0.104 0.003
ID_th 0 0 0 76 19 5 0 0.021 0.149 0.002
ID_ic 0 0 0 99 1 0 0 0.013 0.097 0.004
CPOP 0 0 0 87 12 1 0 0.081 0.122 0.017
NOT (S16) 0 0 1 96 1 2 0 0.015 0.097 0.082
MARS 0 84 13 3 0 0 0 3.720 2.160 0.003
TF 0 0 0 0 0 21 79 0.024 0.433 0.097
TS 0 0 0 96 4 0 0 0.094 0.198 0.110
local_poly 0 100 0 0 0 0 0 0.331 0.956 0.095
DAIS 0 1 2 96 1 0 0 0.461 0.300 0.012
ID_th 0 0 0 92 8 0 0 0.482 0.294 0.006
ID_ic 39 1 15 44 1 0 0 3.100 19.50 0.025
CPOP 0 0 0 89 11 0 0 1.650 0.323 0.025
NOT (S17) 100 0 0 0 0 0 0 6.650 49.00 0.039
MARS 100 0 0 0 0 0 0 6.590 48.00 0.003
TF 0 0 0 0 0 82 18 1.200 0.323 0.110
TS 100 0 0 0 0 0 0 6.650 49.00 0.197
local_poly 100 0 0 0 0 0 0 6.610 5.220 0.639
Appendix C Proof of Theorem 1